# Manuel Morante Kristian Frølich Naveed ur Rehman

### v1

@@ 3.3 Multiscale Functional Connectivity
Overall, Figure 7 indicates similarities among the experiments, particularly within the networks linked to the main modules. A significant activation pattern was observed in modes 2, 3, 4, and 5. On closer inspection, we observed some relevant patterns. First, mode 2 was characterized by extensive activity across several brain networks involving numerous ROIs.
This connectivity suggests a broad engagement of brain regions in this mode, indicating a complex and integrative role in coordinating diverse neural processes. In contrast, Mode 5 demonstrated more focused activation patterns, indicating less extensive networks. This mode’s more localized activity suggests specialized functions, potentially related to specific cognitive tasks. Further visual inspection of the FC matrices showed that the strength and distribution of the correlations within and between brain modules varied significantly across the different modes and experiments.
For instance, modes 3 and 4 exhibited similar connectivity patterns. These modes contributed to overall functional connectivity by integrating and modulating activity across. In mode 2, a notable correlation was observed between the occipital and temporal modules. This connection, however, was considerably weaker in the motor and gambling experiments, suggesting contextual modulation of connectivity. Interestingly, while the temporal cortex and the Default Mode Network (DMN) were disconnected in the resting-state experiment, they were conspicuously linked during task-related activities. This finding highlights brain connectivity dynamics and adaptability across different tasks.
Furthermore, during the gambling experiment in Mode 2, a pronounced left-hemisphere lateralization was linked with the limbic module, showing the potential role of emotional processing in gambling tasks. In contrast, the DMN and the Occipital cortex exhibited a significant correlation during the motor task, which was absent in other experiments. This suggests a specific involvement of these regions in motor control and visual processing.
Mode 3 and 4 mirrored the connectivity patterns of mode 2 but with more intense connections between the occipital and temporal cortex, indicating a potential enhancement or amplification of sensory processing in this mode. Unlike modes 2 and 3, which did not show the same structural pattern, hinting at functional differences even within those modes.
Mode 5 presented a unique scenario; it exhibited almost identical patterns across all experiments but with noticeably reduced connections both within and between modules. Notably, significant activity in the limbic module was observed in its connections with other modules, except in the gambling experiment. This absence might reflect the unique emotional or cognitive demands of the gambling task, which could dissociate limbic activity from other brain regions (Harrison et al., 2021).
In modes 2 and 3, the consistent significant correlation between the occipital and temporal modules indicates strong sensory integration. This is less pronounced in motor and gambling experiments, suggesting a shift of resources towards task-specific networks. Occipital-temporal connectivity is critical for processes such as visual recognition and memory retrieval, which may be differentially recruited depending on the task at hand.
The temporal cortex and DMN’s distinct connectivity in task-related experiments, as compared to the resting state, underlines the task’s influence on the brain’s intrinsic activity. The DMN, typically active during rest and internal mentation, plays an interactive role when engaging with external tasks. This is reflected in enhanced temporal-DMN connectivity.
@@ 4 Limitations and future work
As with any other study, it is important to take into account there are some limitations concerning our study.
Firstly, our analysis was conducted on three fMRI datasets with similar imaging protocols. While these datasets provided valuable insights into MFC behavior and performance, it is important to recognize that the generalizability of our findings may be limited. To fully understand the robustness and applicability of MFC, it would be beneficial to apply this methodology to a broader range of datasets, conditions, and experiments.
Secondly, for simplicity, our study focused on the Region of Interest (ROI) level. This decision was made to simplify the analysis and provide clear, interpretable results. However, the MFC methodology is inherently versatile and can be applied directly at the voxel level as well. Exploring the voxel-level application of MFC could yield fine-grained insights into brain functional connectivity patterns.
Last but not least, in this study, we did not explore in detail the dynamic aspects of the MFC to facilitate the introduction and understanding of the proposed approach. While our study provided preliminary insights, we did not explore completely all the advantages that the temporal time activation patterns of the IMs have to offer. A more thorough examination of temporal activation patterns, connectivity dynamics, and how they relate to conventional dynamic FC approaches is necessary to fully understand MFC capabilities.
Regarding potential paths and future work, there are several promising directions for future research to build upon our current findings and address the identified limitations.
One potential direction is to investigate MFC performance using more advanced algorithms tailored specifically to fMRI data. In our current study, we employed the two most popular algorithms, MEMD and MVMD, with quite different success. We believe that the exploration and development or adaptation of a more sophisticated algorithm could potentially improve the reliability and accuracy of the reported results, leading to a better characterization of the FC.
Another relevant path for future research involves exploring group-level strategies that integrate information from multiple participants. Our current study focused on individual-level analysis for MFC, which exploits MMD’s natural flexibility for accommodating individual variations. Nonetheless, it is still unclear how group-level strategies could be potentially integrated into MFC to accommodate for common information and, potentially enhance the robustness and generalizability of our findings.
@@ 5 Conclusions
In this study, we proposed a novel methodology for extracting neurophysiological functional information from fMRI data across multiple timescales, referred to as Multiscale Functional Connectivity (MFC), by decomposing fMRI activity into distinct intrinsic modes using Multivariate Mode Decomposition (MMD). To the best of our knowledge, this is the first time such an analysis has been performed on fMRI data. Unlike previous similar studies, such as Yue et al. (2019), our proposed method explicitly exploits the multivariate nature of fMRI and accommodates interfering components without manually specified temporal filters. As such, our method provides a more comprehensive view of the underlying fMRI data, which is essential for understanding its dynamics and interactions.
The analysis of three different fMRI experiments revealed that the algorithm Multivariate Variational Mode Decomposition (MVMD) constitutes a suitable candidate for MMD in fMRI data. Our results showed that MVMD provides a meaningful representation consistent with previous research while offering reliable, functional connectivity patterns among participants. The study of functional connectivity patterns among different neurophysiological components sheds light on the intertwined roles of various brain networks at several timescales. We believe that further adoption of MFC would give more insight into fMRI data and enhance our understanding of brain behavior.
Funding
This work has been partly supported by Lundbeck Foundation Investigator network grant R426-2023-133.

### v2

@@ 4.6 Neurophysiological MFC patterns and analysis
From a general overview of Figure 10, we found some similarities among the experiments, particularly within the networks linked to the main modules. We observed some relevant patterns during closer inspection, which we briefly summarize below.
Firstly, the IM 2 was characterized by extensive activity across several brain networks involving numerous ROIs. This connectivity suggests a broad engagement of brain regions in this mode, indicating a complex and integrative role in coordinating diverse neural processes. A notable correlation was observed between the occipital and temporal modules. This connection was considerably weaker in the motor and gambling experiments, suggesting contextual modulation of connectivity. Interestingly, while the temporal cortex and the Default Mode Network (DMN) were disconnected in the resting-state experiment, they were linked during task-related activities. This finding highlights the brain’s connectivity dynamics and adaptability across different tasks.
Furthermore, during the gambling experiment in mode 2 FC patterns, pronounced left-hemisphere lateralization was linked with the limbic module, indicating the potential role of emotional processing in gambling tasks [45]. On the other hand, the DMN and the occipital cortex exhibited a significant correlation during the motor task, which was absent in the different experiments. This suggests a specific involvement of these regions in motor control and visual processing. These observations align with results from FC’s temporal dynamics in task-relevant representations during cognitive control [9].
@@ 5 Limitations and future work
As with any other study, there are some limitations, which we summarized as follows.
  • We analyzed three fMRI datasets with similar imaging protocols. Further experiments with a broader range of datasets and conditions would hepl to fully understand MFC’s robustness and applicability.
  • Exploring dynamic aspects of MFC was beyond the scope of this study. A thorough examination of temporal connectivity dynamics of MFC and how they relate to other dynamic FC approaches could be explored in future work.
  • We focused on a single brain atlas. A comparison with other alternative brain atlas could also provide valuable insights on MFC’s performance.
  • Our study focused on the ROI level. Exploring the voxel-level application of MFC could yield fine-grained insights into the brain FC.
  • The current MMD algorithms are applied individually. More advanced algorithms tailored to fMRI data and group-level strategies that integrate information from multiple participants could be explored in future research.
@@ 6 Conclusions
In this study, we proposed a novel methodology for extracting neurophysiological functional information from fMRI data across multiple timescales, referred to as Multiscale Functional Connectivity (MFC), by decomposing fMRI activity into distinct intrinsic modes using Multivariate Mode Decomposition (MMD). To the best of our knowledge, this is the first time such an analysis has been performed on fMRI data. Unlike previous similar studies, e.g., [23], our proposed approach accounts for the natural nonlinear, nonstationary, and multivariate nature of fMRI while allowing for the particularities of each individual in a data-driven way. As such, MFC provides a more comprehensive view of the underlying fMRI data, which is essential for understanding its dynamics and interactions.
The analysis of three different fMRI experiments revealed that the algorithm Multivariate Variational Mode Decomposition (MVMD) constitutes a suitable candidate for MMD in fMRI data. Our results showed that MVMD provides meaningful results while offering reliable, functional connectivity patterns among participants. The study of these connectivity patterns among different neurophysiological components evidenced the intertwined roles of various brain networks at several timescales. We believe that further adoption of MFC would give more insight into fMRI data and enhance our understanding of brain behavior.